OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. z ÷ 3 – 5z + 4(z – 2) when z = 12 A. –16 B. –96 C. –98 D. –14

OpenStudy (anonymous):

@Godlovesme

OpenStudy (godlovesme):

can u try first @AidanWithanAnota :)

OpenStudy (godlovesme):

like i showed u on the other one

OpenStudy (anonymous):

z / 3 - 5z + 4(z - 2) when z = 12 12/3 - 5(12) + 4(12 - 2) ... can you finish ?

OpenStudy (godlovesme):

yes like @kelliegirl33 did don't forget to do the division and parenthesis first @AidanWithanAnota

OpenStudy (anonymous):

I got 96

OpenStudy (godlovesme):

i got -96

OpenStudy (godlovesme):

4-60+40 ..

OpenStudy (anonymous):

4 - 60 + 40 = -60 + 44 = -16

OpenStudy (anonymous):

Wow I was off quite a bit

OpenStudy (godlovesme):

4-60 = -56 @kelliegirl33

OpenStudy (anonymous):

sometimes it can be a little confusing

OpenStudy (anonymous):

ok..-56 + 40 = -16

OpenStudy (godlovesme):

yes sorry for confusing guys >.< :P

OpenStudy (anonymous):

if the signs are different, subtract and take the sign of the bigger number

OpenStudy (anonymous):

I have 1 more question

OpenStudy (godlovesme):

got it @kelliegirl33 i was rushing sorry @AidanWithanAnota i gtg :( @kelliegirl33 might help